# Boundary of the Moduli Space of Stable Cubic Fivefolds

### v6

@@ 6 Adjacency relations among strictly semistable components
    For each pair of reduced weight vectors $r_{i},r_{j}$ listed in Section 2, we study the slices
    $I(r)_{\geq 0}\cap I(r_{i})_{=0},\qquad I(r)_{\geq 0}\cap I(r_{j})_{=0},$
    and determine whether they are inclusion-maximal subsets inside the corresponding hyperplanes. The same algorithm as in Section 2 applies verbatim.
@@
    If these slices are maximal precisely for $r=r_{j}$ (resp. $r=r_{i}$), then $\Phi_{i}$ and $\Phi_{j}$ admit a common wall degeneration.
@@
    Using the explicit normal forms from Section 4, we verify adjacency by checking that suitable representatives $\widetilde{\varphi}_{i}\in\Phi_{i}$ and $\widetilde{\varphi}_{j}\in\Phi_{j}$ satisfy
    $\lim_{t\to 0}\lambda_{j}(t)\cdot\widetilde{\varphi}_{i}\;=\;\lim_{t\to 0}\lambda_{i}(t)\cdot\widetilde{\varphi}_{j},$
    where $\lambda_{k}$ denotes the 1-PS corresponding to $r_{k}$.

### v7

@@ 8 Non-inclusions and Filters
    Filter 1: Dimension (§8.1). If $\Phi_{k}\subseteq\Phi_{\ell}$ and $k\neq\ell$, then necessarily
    $\dim\Phi_{k}\;<\;\dim\Phi_{\ell}.$
@@
    Filter 2: Apolar Betti numbers (§8.2). Within a fixed Hilbert-function stratum, the graded Betti numbers of the apolar algebra can only increase under specialization; hence an inclusion forces the entrywise Betti-table inequalities
    $\beta_{i,j}\!\bigl(A_{f_{k}}\bigr)\ \geq\ \beta_{i,j}\!\bigl(A_{f_{\ell}}\bigr)\qquad\text{for all }(i,j).$
@@
    Filter 3: Singular-locus Hilbert functions (§8.3). For a cubic $f$, the Hilbert function of the singular scheme $\mathrm{Sing}(X_{f})$ is upper semicontinuous in families, so an inclusion forces the degreewise inequalities
    $\mathrm{HF}_{\mathrm{Sing}(X_{f_{k}})}(d)\ \geq\ \mathrm{HF}_{\mathrm{Sing}(X_{f_{\ell}})}(d)\qquad\text{for all }d\geq 0.$